### v2

# MCA-Bench: A Multimodal Benchmark for Evaluating CAPTCHA Robustness Against VLM-based Attacks

Abstract
As automated attack techniques rapidly advance, CAPTCHAs remain a critical defense mechanism against malicious bots. However, existing CAPTCHA schemes encompass a diverse range of modalities—from static distorted text and obfuscated images to interactive clicks, sliding puzzles, and logic-based questions—yet the community still lacks a unified, large-scale, multimodal benchmark to rigorously evaluate their security robustness. To address this gap, we introduce MCA-Bench, a comprehensive and reproducible benchmarking suite that integrates heterogeneous CAPTCHA types into a single evaluation protocol. Leveraging a shared vision–language model backbone, we fine-tune specialized cracking agents for each CAPTCHA category, enabling consistent, cross-modal assessments. Extensive experiments reveal that MCA-Bench effectively maps the vulnerability spectrum of modern CAPTCHA designs under varied attack settings, and—crucially—offers the first quantitative analysis of how challenge complexity, interaction depth, and model solvability interrelate. Based on these findings, we propose three actionable design principles and identify key open challenges, laying the groundwork for systematic CAPTCHA hardening, fair benchmarking, and broader community collaboration. Datasets are available at https://www.kaggle.com/datasets/luffy798/mca-benchmultimodal-captchas. Code is released at https://github.com/noheadwuzonglin/MCA-Bench.
Figure 1: Data samples from MCA-Bench includes four categories and 20 sub-clusters of Point-and-Click Localization, Static Visual Recognition, Textual Logic Q&A and Interactive Manipulation.
1 Introduction
Table 1: Overview of CAPTCHA datasets for various task types. The dataset includes multiple task categories, such as static visual recognition, point-and-click localization, interactive manipulation, and textual logic Q&A tasks, with varying image sizes and training samples to support diverse CAPTCHA research.
Type Dataset Image size #Images #Train Original image Single image
Static visual recognition Classic character CAPTCHA 160×60 5000 4800
Distorted-word CAPTCHA 800×200 5000 4800
Arithmetic-character CAPTCHA 130×48 5000 4800
Point-and-click localization 3×3 grid selection 672×672 5500 5300
Inverted-letter selection 224×224 5000 4800
Geometric-shape recognition 224×224 5898 5698
Brightness discrimination 224×224 5377 5177
Hollow-pattern recognition 224×224 5000 4800
Sequential-letter ordering 300×150 4999 4799
Full-image grid selection 512×512 5015 4815
Color discrimination 224×224 5000 4800
Vowel selection 224×224 5322 5122
Arithmetic-selection 240×240 5065 4865
Rotated-letter selection 224×224 5015 4815
3×3 jigsaw-swap selection 224×224 6505 6305
Interactive manipulation Sliding-block 600×360 5012 4812
Rotation-block 600×360 5041 4841
Alignment sliders 600×360 5006 4806
Textual logic Q&A Text-based arithmetic 1000 800
Commonsense reasoning 1000 800
CAPTCHA (Completely Automated Public Turing test to tell Computers and Humans Apart) has long played a vital role in protecting online services from automated attacks. In recent years, growing research interest in AI security topics such as adversarial attacks van2023anti ; song2024anti ; song2025idprotector and digital watermarking ci2024ringid ; ci2024wmadapter ; liu2024image ; yang2024can has also drawn attention to the vulnerability of CAPTCHA systems. However, with rapid advances in deep learning, computer vision, and multimodal pretraining models, many CAPTCHA types once considered secure have now become susceptible to machine-learning-based attacksbursztein2011text ; sivakorn2016robot ; ye2018yet . Techniques such as GANs, vision-language models (VLMs), and reinforcement learning have enabled attackers to mimic human behavior with increasing precisiongoodfellow2014generative ; radford2021learning ; li2022blip ; shah2023lm . As a result, researchers have begun developing multimodal CAPTCHA datasets and evaluation frameworks to assess model performance across various CAPTCHA typesacien2020becaptcha ; acien2021becaptcha ; gao2021research . This makes it vital to reassess CAPTCHA’s real-world security to ensure a trustworthy Internet service ecosystemnoury2020deep ; gao2021research .
Despite the emergence of new interactive CAPTCHA forms, existing studies often focus on specific categories (e.g., text recognition, image selection) without offering comprehensive comparisonsbursztein2011text ; ding2025illusioncaptcha ; hernandez2021basecass . Moreover, there is no large-scale, multimodal benchmark for evaluating CAPTCHA designs against intelligent-agent attackshossen2022aaecaptcha ; searles2023empirical ; deng2024oedipus . This gap limits the identification of vulnerabilities and the development of more robust CAPTCHA systems. Therefore, a unified, standardized evaluation platform is urgently needed for security analysis across modalities and tasks.
MCA-Bench is the first end-to-end CAPTCHA security benchmark spanning four modalities—static visual recognition, point-and-click localization, interactive manipulation and textual logic Q&A—across twenty real-world tasks. It provides over 180000 training samples and a 4000-item test set (200 examples per task), organized into four clusters that respectively evaluate OCR robustness to visual noise, target retrieval in complex scenes, human-like interaction behaviors, and multi-step language reasoning. Representative samples from the MCA-Bench dataset are shown in Figure 1.
We use QwenVL-2.5-7B as the vision-language backbone, fine-tuned with LoRA adapters for each task. Training for static and logic CAPTCHAs is supervised with target labels, while for interactive tasks, human demonstration data is used for behavior cloning. A specially designed JSON protocol facilitates large-scale evaluation and integration.
The evaluation framework uses pass rates as the core metric. Experiments show that multimodal VLMs achieve over 96% accuracy on simple tasks, while more complex ones (requiring physical interaction or multi-step logic) still pose significant challenges, with success rates as low as 2.5%. This highlights the coupling of ’visual confusion, depth of interaction, and semantic complexity’ with attack difficulty, providing actionable insights for future CAPTCHA design. MCA-Bench is released open-source, enabling easy replication and fostering ongoing cycles of attack, defense, and benchmark iteration. The main contributions are as follows.
  • MCA-Bench: the first large-scale, cross-modal CAPTCHA attack benchmark with 20 real tasks.
  • Proposed a unified evaluation pipeline with a single pass-rate metric and open source scripts.
  • First full-scale CAPTCHA security assessment with guidance for human–machine verification.
Table 2: Comparison of pass rates for various CAPTCHA types. This table presents the pass rates for different CAPTCHA task types, comparing the performance of the QwenVL-2.5-7B-Instruct model with human pass rates.
CAPTCHA Type QwenVL-2.5-7B-Instruct Pass Rate Human Pass Rate
Point-and-Click Localization 3×3 grid selection 0.96 0.88
Inverted-letter selection 0.52 0.94
Geometric-shape recognition 0.96 0.98
Brightness discrimination 0.665 0.78
Hollow-pattern recognition 0.995 0.98
Sequential-letter ordering 0.925 0.98
Full-image grid selection 0.35 0.74
Color discrimination 0.99 0.88
Vowel selection 0.975 0.80
Arithmetic selection 0.025 0.78
Rotated-letter selection 0.335 0.74
3×3 jigsaw-swap selection 0.96 0.80
Static Visual Recognition Classic character CAPTCHA 0.32 0.92
Distorted-word CAPTCHA 0.985 0.84
Arithmetic-character CAPTCHA 0.695 0.98
Textual Logic Q&A Text-based arithmetic 0.985 0.97
Commonsense reasoning 0.97 0.86
Interactive Manipulation Sliding-block 0.36 0.74
Rotation-block 0.28 0.76
Alignment sliders 0.55 0.72
2 Related Work
2.1 VLMs in Structured Image Tasks
Visual Language Models (VLMs) bridge images and text through unified embeddings, with early methods like CLIP radford2021learning and ALIGN pmlr-v139-jia21b excelling in simple visual tasks but falling short in complex reasoning. Recent models integrate large language models (LLMs) with visual encoders to enhance cross-modal understanding—BLIP-2 li2022blip uses frozen components for efficient training, while MiniGPT-4 DBLP:conf/iclr/Zhu0SLE24 maps visual features to LLMs for stronger image-text QA. As research shifts toward structured tasks like tables and documents, which resemble CAPTCHA images with dense text and noise, models such as Qwen-VL wang2024qwen2vlenhancingvisionlanguagemodels improve encoder design and pretraining for better layout and multilingual comprehension. Architectures like LLaVA NEURIPS2023_6dcf277e and MiniGPT-4 DBLP:conf/iclr/Zhu0SLE24 adopt a unified visual projection and instruction-tuning framework, achieving robustness on distorted and occluded inputs. However, most VLMs lack CAPTCHA-specific adaptation, limiting their accuracy and generalization.
2.2 Evolution of Intelligent Agents Toward AGI
An agent in AI refers to an autonomous entity that perceives its environment via sensors, makes decisions, and acts through actuators russell2021modern , characterized by autonomy, reactivity, and social interactivity—key traits driving progress toward Artificial General Intelligence (AGI) 10.1007/3-540-58095-6_9 ; 109c6f85f2734b649d5f7b7a1946b2c1 . Agent research has evolved from early symbolic agents with interpretable rule-based reasoning 10.1145/360018.360022 ; 10.5555/2974989 , yet limited in handling uncertainty and scalability 10.5555/110787 , to reactive agents offering fast perception-action loops 10.5555/1625995.1626091 , albeit with limited decision-making. The rise of reinforcement learning (RL) enabled agents like AlphaGo 44806 to learn from interactions, though challenges like sample inefficiency and poor generalization remained 10.1023/A:1015008417172 ; 1996cs……..5103K ; 1571980075547310080 ; 1573668924277769344 . Transfer learning and meta-learning improved task adaptation via knowledge reuse and learning-to-learn paradigms 10.5555/2772879.2772905 ; 2015arXiv151106342P ; 10.1109/TPAMI.2023.3292075 ; 2016arXiv161102779D ; 10.5555/3305381.3305498 ; 10.5555/3327345.3327436 ; rakelly*2019efficient ; Fakoor2020 , despite issues in task diversity and pre-training costs farebrother2019generalization . Recently, LLM-based agents have transformed the field, exhibiting emergent reasoning and planning abilitiesmckinney2020training ; achiam2023gpt ; wei2022emergent ; mann2020language , multimodal understanding (e.g., BLIP-2 li2023blip ), dynamic task decomposition (e.g., Voyager wang2023voyager ), and tool usage (e.g., Toolformer schick2023toolformer ), enabling general-purpose intelligence park2023generative ; liu2023training ; sumers2024cognitivearchitectureslanguageagents ; Wang_2024 . Their capabilities in zero-shot generalization wei2021finetuned ; sanh2021multitask ; chung2022scalinginstructionfinetunedlanguagemodels and social collaborationhong2024metagptmetaprogrammingmultiagent mark a paradigm shift in agent research.
2.3 Advances and Challenges in CAPTCHA Security
Early text CAPTCHAs used heavy distortion and noise, but segmentation flaws and CNN-based attacks soon achieved high recognition rates—even breaking reCAPTCHA under laboratory conditions chellapilla2005building ; simard2003best ; bursztein2011text . Image-classification schemes such as ASIRRA shifted effort to visual cognition, yet SVMs trained on public datasets ultimately compromised them elson2007asirra ; golle2008machine . Contemporary defences leverage deep object-detection networks alongside generative techniques (e.g., diffusion models and style-transfer) to create more diverse and intricate puzzles nian2022deep ; xing2023review ; jiang2023diff ; ding2025illusioncaptcha ; goodfellow2014generative ; luo2024image . Interactive CAPTCHAs, exemplified by reCAPTCHA v2, further blend visual tasks with user gestures such as drag-and-drop or click behaviour to raise the bar for automation shet2014you . Despite these advances, CAPTCHA schemes still struggle with adaptive difficulty tuning, behaviour consistency across devices, adversarial robustness, and—critically—the absence of standardized, reproducible benchmarks for end-to-end evaluationkumar2021benchmarks ; alsuhibany2016benchmark .
Figure 2: Schematic overview of the MCA-Bench data-acquisition and annotation workflow. The four grey panels—from left to right—represent the dataset categories Static Visual Recognition, Interactive Manipulation, Point-and-Click Localization, and Textual Logic Q&A. The red labels (SVT, MIM-RU, CPTL, and LLM-GFP) denote the dedicated data-collection pipelines associated with each category. Each pipeline follows a unified four-stage paradigm: (i) determining the raw input format; (ii) applying task-specific geometric transformations, coordinate projections, or prompt-template generation; (iii) separating fine-grained annotation types; and (iv) saving the annotations to text files.
3 Dataset Construction Process
This section outlines the MCA Bench pipeline—from raw sample collection to final benchmark release—while detailing the data-processing and annotation strategies for its four task clusters and twenty subtasks. The complete workflow appears in Figure 2.
3.1 Data Collection Sources and Compliance Strategies
In developing the MCA-Bench benchmark suite, we implemented a comprehensive data collection and compliance management system to uphold scientific integrity and ensure legal and ethical standards in evaluating CAPTCHA security. Our approach combines independently created data and reused public datasets. For the former, we designed a diverse CAPTCHA dataset, including distorted text, obfuscated graphics, interactive tasks, and puzzles, based on a thorough analysis of technology trends and security needs. These designs are original, validated, and proprietary to our team.
For reused datasets, we carefully selected authorized academic, open-source, and industry resources, adhering to licensing agreements and privacy regulations. We ensured compliance through preprocessing steps like anonymization and mitigating privacy risks. Additionally, we established a multi-level review system with legal and ethics oversight, encrypted storage, and tiered access control to guarantee ongoing compliance and data security for MCA-Bench.
3.2 Data Collection and Processing
3.2.1 Text-based Task Data Collection
For text-based CAPTCHA tasks, we developed a semi-automated LLM-driven pipeline to efficiently generate and filter math and commonsense questions, reducing manual effort while maintaining quality and diversity. By using adaptive prompts, we guided the Qwen LLM to create structured, semantically relevant question-answer pairs, incorporating knowledge constraints, task-guided sampling, and type-control parameters for difficulty and logic. The output was further refined through manual filtering with a custom evaluation protocol, assessing grammar, reasoning, and ambiguity to ensure robustness and clarity for real-world use.
3.2.2 Click-based Coordinate Task Data Collection
We propose CPTL (Click-based Positioning and Target Localization), a multimodal benchmark designed to evaluate models’ spatial localization accuracy and image-language instruction alignment across varying scenario complexities. The framework consists of two main tasks: Free-form Pointing and Grid-based Discrete Pointing. In the Free-form Pointing task, we use a highly controllable image generation process combining procedural background perturbations with public visual datasets, creating images layered with semantic content, based on Flickr scenic photos. The Grid-based Discrete Pointing task employs a standardized 3×3 grid layout to assess model decision-making within constrained regions, dividing each image into 9 segments (0 to 8) to form a discrete click space. The image content is derived from a 64-class animal classification datasetanthony_therrien_2024 , Flickr scenic backgrounds, Google ReCaptcha V2 challenges, and manually designed targets.
CPTL supports systematic, adjustable evaluation of spatial complexity and instruction modalities, enabling isolated assessment of spatial parsing and language comprehension. Its key advantage lies in simulating real-world CAPTCHA click scenarios while maintaining robust model performance under noisy backgrounds, multi-target interference, and weak semantic prompts. Experimental results demonstrate that multimodal models fine-tuned with CPTL exhibit strong localization and semantic understanding, providing a reliable evaluation baseline and optimization direction for future CAPTCHA security research.
3.2.3 Static Visual Recognition Task Data Collection
We propose SVT (Static Visual Textual Understanding), a framework designed to evaluate multimodal models’ ability to understand and reconstruct text from distorted images. SVT combines procedural image generation with language modeling constraints, allowing high controllability and complexity. It uses character-level geometric perturbations (e.g., affine transformations, warping, and perspective distortion) to generate challenging yet recognizable character patterns, while a stochastic interference generator adds random noise and occlusions. SVT effectively tests models’ character-level attention and resilience to structural noise, providing a comprehensive evaluation from visual perception to symbolic reasoning. Results show that SVT reveals bottlenecks in fine-grained character perception and serves as a valuable benchmark for assessing the robustness of pretrained multimodal models.
3.2.4 Interactive Behavior Task Data Collection
We present a multimodal interaction modeling framework for interactive CAPTCHA tasks, based on real user operations, to enhance vision-language models’ dynamic understanding capabilities. Focusing on continuous motion interactions—such as sliding alignment, rotation calibration, and trajectory restoration—the framework requires models to perceive spatial structures, motion directions, and behavioral patterns. Data is collected through large-scale manual sampling across devices, capturing thousands of diverse user trajectories with timestamps and velocity, enabling realistic spatiotemporal behavior modeling.
Tasks are designed around target state restoration, formalized into standardized sub-tasks with defined start, goal, and intermediate states. Interaction trajectories are serialized into behavioral vector streams and jointly encoded with image content to guide visual-behavior learning. Fine-tuning with high-quality samples improves the model’s decision consistency and action alignment. To evaluate performance, we introduce metrics including Center Deviation Error, Angular Restoration Accuracy, Slide Path Alignment Rate, and Motion Variability Index, offering comprehensive assessment beyond traditional image recognition benchmarks.
3.3 Data Annotation Strategy
To ensure consistent training and evaluation of multimodal CAPTCHA tasks, we developed a task-driven annotation framework with standardized labeling for four representative task types.
3.3.1 Text-based Tasks: Semantic Consistency Annotation Based on Q&A Structure
We propose a semi-automated, LLM-driven pipeline for generating and filtering math and commonsense questions in text-based CAPTCHA tasks, significantly reducing manual workload while maintaining question quality and diversity. An adaptive prompt template guides the Qwen LLM to produce well-structured and semantically coherent question–answer pairs. Unlike traditional approaches relying on random sampling or manual composition, our method integrates knowledge-constrained generation and task-guided sampling, with type-control parameters to regulate structural complexity, difficulty, and logical patterns.
Following generation, a manual filtering stage applies a multi-dimensional evaluation protocol—assessing grammatical correctness, reasoning consistency, and task ambiguity—to ensure that questions are both discriminative and robust for real-world CAPTCHA evaluation.
3.3.2 Unified Intent Modeling for Coordinate Pointing and Grid Selection
For free-form coordinate pointing tasks, we use a box-to-center projection strategy. Annotators mark the top-left and bottom-right corners of the target’s bounding box using absolute pixel coordinates, with (0, 0) at the top-left of the image. The geometric center of each box serves as the training supervision signal. During inference, a prediction is considered correct if it falls within the corresponding bounding box, following an IoB-Gated Validation mechanism. This approach offers spatial tolerance, improves robustness to outliers, and enhances training convergence.
Formally, the validation criterion between the predicted point $p$ and the bounding box $b=[\,b_{\min},\,b_{\max}\,]$ is defined as:
$\mathcal{G}\bigl{(}p,\,b\bigr{)}\;=\;\mathbb{I}\Bigl{(}\bigl{\lVert}\,D^{-1}\bigl{(}p-\tfrac{1}{2}(b_{\min}+b_{\max})\bigr{)}\bigr{\rVert}_{\infty}\leq\tfrac{1}{2}\Bigr{)}$ (1)
Here, $\mathbb{I}(\cdot)$ is the indicator function (1 if the condition holds, 0 otherwise), $D=\operatorname{diag}(\,{b}_{\max}-\,{b}_{\min})$ is the diagonal matrix of bounding-box width and height, and $\lVert\cdot\rVert_{\infty}$ is the $\ell_{\infty}$ norm (the maximum absolute component). Thus, the expression tests whether the predicted point lies within the normalized box region centered on the ground-truth box.
We adopt a Discrete Region Index Supervision scheme for grid-based pointing, where each image is divided into a uniform 3×3 grid, yielding nine equal-sized cells indexed from 0 to 8 in row-major order. Annotators label the target region by selecting the corresponding grid index, which serves as a discrete classification target during training. This Semantic-Aligned Index Encoding effectively aligns linguistic cues with local image regions, simplifies supervision under multi-object interference, and enhances model robustness and generalization in the face of linguistic ambiguity and similar visual distractors.
3.3.3 Static Visual Recognition Class Task Labeling Methods
We developed a refined labeling system for static visual recognition tasks to address task diversity and challenges. For character recognition, we use character-level text tagging, where each character in the image is directly labeled and aligned with the content for accurate learning. Labels are automatically generated and validated to prevent errors, with random sampling and manual review to ensure readability and consistency.
For semantic tasks, such as mathematical expressions or symbolic combinations, we apply Visual-Symbolic Abstraction Parsing, parsing the image’s semantic structure into a "final semantic solution". The labels are deterministic and unique, supporting logical reasoning and computation. This system ensures clear, consistent, and task-aligned supervision, providing precise signals for multimodal models and reliable evaluation in complex scenarios.
3.3.4 Interactive Behavior Task Labeling Methods
The annotation system for interaction-based CAPTCHAs focuses on modeling user behavior through detailed trajectory data, target operations, and motion features across common interaction modes such as sliding, rotation, and layer alignment. These tasks combine visual restoration (e.g., aligning puzzle pieces or rotating images) with behavioral trajectory monitoring, allowing the system to distinguish between natural user actions and scripted or emulated inputs. Real user behaviors show non-linearity, irregular speed, and realistic timing, unlike the overly uniform patterns of bots or automation tools.
In slider CAPTCHAs, users drag a puzzle piece into a matching gap; in rotation-based ones, they adjust the image angle via a slider; and in alignment CAPTCHAs, they slide misaligned layers back into position. During these actions, 10 monitoring points capture behavioral data including position, time, and event type. These are analyzed for trajectory naturalness, speed/acceleration profiles, and duration plausibility (data fields detailed in Appendix). For rotation tasks, there is a mapping between the slider’s horizontal displacement ($\Delta x$) and the counter-clockwise rotation angle ($\phi$) of the image. This mapping incorporates affine transformations and periodic functions. Given a maximum slider travel distance of $L=250$ pixels, the rotation angle $\phi$ can be computed as follows:
$\phi(\Delta x)=\arg\left(\exp\left(i\cdot 2\pi\cdot\frac{\Delta x}{L}\right)\right)\cdot\frac{360^{\circ}}{2\pi}$ (2)
[…]
Figure 3: The schematic diagram illustrates the data flow and key module configuration across the four stages of the end-to-end framework: unified interface access, gent fine-tuning loading, collaborative inference execution, structured result feedback.
4 Agent Pipeline and Experiments
This section presents a detailed analysis of the experimental results based on the MCA-Bench benchmark.
Set up We use Qwen2.5-VL-7B-Instruct as the backbone and fine-tune LoRA adapters on 20 CAPTCHA tasks. Images are resized to 224×224, with inputs paired to structured prompts and outputs. Training runs on 4 H20 GPUs with per-GPU batch size 8 and gradient accumulation of 4 (effective batch size 32). We optimize using AdamW with a linearly decaying learning rate starting at $1\times 10^{-4}$. Checkpoints are saved every 100 steps, with early stopping triggered if validation loss fails to improve over 20 evaluation intervals.
MCA-Bench is a high-quality dataset designed to enhance a model’s cracking capabilities We fine-tuned the QWen-2.5VL-7B-instruct model using MCA-Bench, significantly improving its performance. As shown in Figure 4, the fine-tuned model outperforms both the zero-shot baseline and the best closed-source model, ChatGPT-4o, on various CAPTCHA tasks. The dataset boosts the model’s adaptability in visual recognition, logical reasoning, and interaction tasks, while also improving its robustness against complex CAPTCHAs, proving its value as a powerful resource for improving model performance.
Figure 4: Violin plots of Passing-Rate (0–1) for Original Qwen-2.5 VL-7B-Instruct, Original ChatGPT-4-o, and Finetuned Qwen-2.5 VL-7B-Instruct on 20 CAPTCHA classes. Outlines show density; boxes, dots, and whiskers mark IQR, median, and range.
MCA-Bench Dataset Overview MCA-Bench is the first multimodal dataset offering a wide range of CAPTCHA tasks for research. It covers visual recognition, point selection, textual reasoning, and interactive operations, with samples of varying image sizes and features. The training set includes a large number of samples, supporting large-scale training and improving model generalization. The dataset also allows multi-image fine-tuning, optimizing the model’s performance across related tasks. Some tasks use publicly available resources, ensuring diversity and reproducibility in research. Table 1 summarizes key details of the dataset, including categories, image sizes, and training quantities.
Agent System Pipeline The system uses an end-to-end pipeline to unify images, text prompts, and user interactions into a single JSON payload. This payload is directly mapped to a semantic space without intermediate parsing. Based on the task ID in the JSON header, the pipeline activates the appropriate LoRA-Agent, each applying fine-tuned parameters to a shared frozen backbone for efficient multi-task operation with low memory overhead. During inference, visual and textual embeddings are routed to the corresponding agent, which produces outputs such as coordinates, text, mouse movements, or character recognition. These are reassembled into the original JSON format for downstream processing, enabling flexible handling of diverse CAPTCHA tasks, and the overall fine-tuning flow is shown in Figure 3.
Fine-Tuned Model Performance in Visual Tasks For simple visual tasks with minimal background interference, the fine-tuned visual-language model outperforms humans in both speed and accuracy, especially in tasks like target recognition and geometric shape identification. However, as task complexity increases—such as with distorted or rotated images—the model’s performance drops significantly, while humans maintain strong performance in these challenging scenarios. This highlights the limitations of visual-language models in handling complex visual transformations.
Human vs. Model Performance in Reasoning and Interaction Tasks In tasks requiring reasoning or behavioral interaction, the fine-tuned model still lags behind humans. Interactive tasks, such as sliding puzzles and rotating tiles, require precision and coordination, which are challenging for the model. Similarly, in tasks demanding multi-step reasoning or common-sense judgment, the model struggles with complex tasks that involve deeper understanding and contextual reasoning. As shown in Table 2, despite the fine-tuned Qwen-Agent’s success in some standard visual recognition tasks, humans still outperform the model in complex tasks involving reasoning and interaction.
5 Design Principles for Next-Generation Cognitive-Secure CAPTCHAs
Large-scale experiments on MCA-Bench demonstrate that fine-tuned multimodal models achieve over 96% success on visual or shallow-semantic CAPTCHAs but drop below 2.5% for tasks requiring physical interaction or multi-step reasoning. This emphasizes the need for a more robust verification approach, as single-dimension obfuscation is no longer effective against vision–language model-based attacks. To address this, we propose a framework that tightly integrates perception, action, and reasoning through three core principles: Deep Modality Coupling, Behavior-Anchored Validation, and Session-Specific Semantic Personalization.
Deep Modality Coupling involves a continuous challenge pipeline that blends visual cue localization, logical inference, and interactive input. This approach dynamically adjusts task complexity based on user performance, enhancing both security and user experience. Behavior-Anchored Validation analyzes user interaction trajectories, granting verification based on similarity to genuine human data, which prevents scripted attacks without hindering user flow. Session-Specific Semantic Personalization embeds a unique semantic seed into each CAPTCHA, ensuring that solutions rely on session-specific context, preventing automated solvers from precomputing answers. Together, these principles create a cognitively rich verification system that is resilient to advanced attacks.
Conclusion
We present MCA-Bench, the first unified benchmark for evaluating CAPTCHA robustness across twenty real-world tasks spanning visual recognition, text comprehension, logical reasoning, and interactive manipulation. Through extensive experiments with LoRA-based attack agents on state-of-the-art vision–language models, we identify structural vulnerabilities in visual and shallow semantic CAPTCHAs, while tasks involving physical interaction and multi-step reasoning demonstrate greater resilience. Based on these findings, we propose three practical design principles—Deep Modality Coupling, Behavior-Anchored Validation, and Session-Specific Semantic Personalization—to guide the development of more secure human–machine verification systems. MCA-Bench establishes a standardized foundation for adversarial evaluation and future research into adaptive difficulty, human behavior modeling, and usability-aware CAPTCHA design.
Supplementary Material
Outline
A More Details on MCA-Bench A
A.1 Overview of Data Sets A.1
A.2 Distribution Pattern and Training Balance Details for Each Type of Task A.2
A.3 Image Resolution Hierarchy A.3
B Market Research on Mainstream CAPTCHA Datasets B
B.1 Comparative Analysis of User Experience and Security Performance B.1
B.2 Evaluation of Mainstream CAPTCHA Types B.2
B.3 Application Trends and User Acceptance Analysis of Mainstream CAPTCHA Types B.3
B.4 Evaluation and Analysis of Mainstream CAPTCHA Types Selection B.4
B.5 Analysis of CAPTCHA Error-Proneness and User Self-Recovery B.5
B.6 Semiotic Analysis of CAPTCHA Types B.6
C More Extensive Experiments C
C.1 Comparison of Model Stability and Human Performance Variability C.1
C.2 Error-Correction Capability Analysis C.2
C.3 Comparative Analysis of Micro-Decision Path Interpretability C.3
D Limitations and Future Work D
D.1 Limitations D.1
D.2 Future Work Outlook D.2
Appendix A More Details on MCA-Bench
A.1 Overview of Data Sets
To systematically compare the adaptability and data support of different CAPTCHA tasks in model training, this study categorizes them into four major types: static visual recognition, point-based localization, incremental interaction, and textual logical reasoning. We summarize seven key technical indicators for representative datasets, including image resolution, data volume, training set size, original annotation attributes, and whether single-image tasks are supported. Character-based CAPTCHAs typically feature large data volumes and low resolutions, providing abundant and low-cost training samples for deep learning models. In contrast, point-click and interaction-based tasks have relatively smaller datasets but higher annotation complexity, often utilizing high-resolution images that significantly enhance models’ fine-grained representation capabilities in multimodal reasoning tasks such as object localization and action prediction. Interactive CAPTCHAs rely on dynamic generation mechanisms, simulating temporal dependencies and adversarial human-machine interactions in real-world scenarios. For these interactive CAPTCHAs, we have specifically designed a sampling field, as detailed in Table 3.
Table 3: Field specification of the CAPTCHA interaction dataset. We define 13 structured fields used to record user interactions with CAPTCHA tasks, including spatial position, timing, image paths, and movement characteristics. These fields are essential for modeling behavior patterns and evaluating automated solver performance.
Name Type Length Decimal point Not null Virtual Key Explanatory note
id smallint 5 1 Primary key ID (maximum 50,000)
data_id int 10 Group ID to which the data belongs ($\leq$10,000)
group_index int 10 Group number
type varchar 10 CAPTCHA type
bg_img_path varchar 50 Background image path
path varchar 50 Template image path (optional)
event_type varchar 10 Event type
x smallint 5 X coordinate ($\leq$10,000 px)
y smallint 5 Y coordinate ($\leq$10,000 px)
t smallint 5 Time value ($\leq$10,000 ms)
speed decimal 10 5 Speed (px/ms), keep 5 decimal places
acceleration decimal 10 5 Acceleration (px/ms2), keep 5 decimal places
jitter_frequency decimal 10 5 Jitter frequency (times/ms), keep 5 decimal places
A.2 Distribution Pattern and Training Balance Details for Each Type of Task
To ensure balanced model performance across diverse tasks, we conducted a visual analysis of the overall dataset structure by subtask type, as presented in Appendix A. As illustrated in Figure 5, the sample distribution across the four task categories—static visual recognition, point-based localization, interactive operation, and textual logical reasoning—remains relatively balanced. This distribution is intended to prevent any single task from dominating the training process, thereby enabling a more comprehensive multimodal learning environment. Moreover, the comparable level of sample support across tasks provides a solid data foundation for subsequent task switching and joint optimization within the model.
Figure 5: Category distribution of CAPTCHA types in the dataset The dataset covers a wide range of 3D-interactive, text-based, and visually complex CAPTCHA categories, with each type contributing approximately 5% of the total. The lowest-frequency categories (e.g., commonsense reasoning and text-based arithmetic) represent specialized reasoning-based challenges, while the most common types focus on perceptual and motor interactions.
A.3 Image Resolution Hierarchy
In terms of model input preprocessing strategies, Figure 6 illustrates the distribution of image resolutions across the entire dataset, serving as an important reference for multi-scale feature learning and computational resource allocation. As shown in Figure 6, the majority of image samples fall within a medium-resolution range, while instances with extreme resolutions are relatively rare. This visualization supports the effectiveness of using a unified resizing and cropping approach to accommodate the detail requirements of most samples efficiently. Additionally, by incorporating multi-resolution data augmentation strategies, model robustness to varying image scales can be further enhanced without significantly increasing computational overhead.
Figure 6: Distribution of image file sizes in the dataset Over 60% of the images are smaller than 100 kB, indicating a strong skew toward low-resolution or compressed images. The dashed line shows the estimated probability density function. Percentage labels denote the relative frequency per bin.
Appendix B Market Research on Mainstream CAPTCHA Datasets
We conducted a comprehensive survey of the mainstream CAPTCHAs currently available on the market, which provided clear requirements and technical references for the development of MCA-Bench.
B.1 Comparative Analysis of User Experience and Security Performance
To systematically evaluate the performance of various mainstream CAPTCHAs in real-world scenarios, this study conducts a comparative analysis across five dimensions: user completion time, misclick rate, visual fatigue, cracking cost, and typical application scenarios, as summarized in Table 4. The results reveal significant differences among CAPTCHA types in terms of usability and security. Color recognition and slider puzzles offer the fastest operation and lowest visual load, providing a superior user experience overall. In contrast, math-based CAPTCHAs require additional cognitive processing and thus result in the longest completion times, though they have relatively low misclick rates. Traditional text-based CAPTCHAs are cognitively demanding, error-prone, and increasingly vulnerable due to advances in automated attack techniques.
In terms of security robustness, slider puzzles are the most resilient, owing to their reliance on authentic user-generated trajectories, making them difficult to bypass using automated scripts. Conversely, traditional text and math CAPTCHAs are more susceptible to automated cracking. Regarding application scenarios, slider puzzles are best suited for high-security contexts such as financial logins; color recognition CAPTCHAs are ideal for systems prioritizing usability or targeting children and elderly users; and traditional text CAPTCHAs are only recommended for low-security environments where legacy compatibility is a concern.
Table 4: User experience and security performance across mainstream CAPTCHA types. This table compares various CAPTCHA mechanisms based on usability and security dimensions, highlighting trade-offs in effectiveness, user preference, and resistance to automation.
CAPTCHA Type Completion Time (s) Mistouch Rate Fatigue (1–5) Cracking Cost (%) Applicable Scenarios
Slider Puzzle CAPTCHA 5–8 15% 2 80% Login for financial apps, enterprise operations
Shape Recognition CAPTCHA 7–12 18% 3 60% Design platforms, graphical UIs
Nine-grid Click CAPTCHA 6–10 20% 3 40% E-commerce registration, social logins
Mathematical Calculation CAPTCHA 10–15 10% 3 20% Educational systems, bank transfers
Color Recognition CAPTCHA 4–7 12% 2 40% Children’s apps, elderly systems
Distorted Character CAPTCHA 7–10 18% 3 20% Old systems, non-core entrances
Traditional Character CAPTCHA 6–9 25% 4 20% Test pages, low-priority verifications
Table 5: General public perspectives on commonly used CAPTCHA systems. We present a comparative analysis of user preferences, perceived security rankings, and subjective commentary reflecting real-world usability and acceptance.
CAPTCHA type Security strength User preference Standpoint
Behavioral CAPTCHA NO.1 $\checkmark$ The operation of sliding is intuitive, and it combines interactivity with relatively high security. Its dynamic feature makes it difficult for automated bots to precisely imitate human behavior
Dynamic Graphic CAPTCHA NO.2 $\times$ By presenting changing visual elements, it increases the difficulty of recognition. However, it may cause visual fatigue, and advanced image analysis algorithms may be able to decipher its patterns over time
Shape Judgment CAPTCHA NO.3 $\checkmark$ It requires users to judge shapes according to certain rules. This type of CAPTCHA makes use of human spatial perception and shape recognition abilities. Although it has a certain security level, sophisticated shape analysis algorithms may pose a threat to its security
Nine-grid Click CAPTCHA NO.4 $\times$ It presents a nine-grid layout with elements and requires users to click on elements that meet specific rules. It balances security and user-friendliness, but if the rules are repetitive, automated scripts may be able to bypass it
Calculation Problem CAPTCHA NO.5 $\checkmark$ It presents arithmetic or logical problems and requires users to calculate and input answers. It tests human cognitive abilities, but simple calculation patterns can be easily automated, reducing its security over time
Color Recognition CAPTCHA NO.6 $\checkmark$ It asks users to recognize colors or select elements based on color-related rules. It is relatively easy for humans but can be a challenge for bots. However, advanced color detection algorithms may undermine its security
Distorted Character CAPTCHA NO.7 $\times$ It displays characters in distorted forms, increasing the difficulty of optical character recognition. However, with the continuous progress of OCR technology and machine learning, its security is lower compared to more interactive types
Traditional Character CAPTCHA NO.8 $\times$ Composed of standard characters, it can be easily recognized by modern OCR software. It offers minimal security and has become obsolete due to its simplicity and vulnerability
B.2 Evaluation of Mainstream CAPTCHA Types
To comprehensively understand the trade-offs between security and user experience in mainstream CAPTCHA types, this study systematically reviews various commonly used CAPTCHA forms in the market, summarizing their technical characteristics and user perspectives, as detailed in Table 5. Analysis across three key dimensions reveals a clear trade-off between security and usability: behavioral CAPTCHAs, which leverage sliding trajectories and temporal patterns, offer the highest level of security. They are followed by dynamic graphics and shape-based CAPTCHAs, which rely on motion interference or spatial recognition. However, as the level of interaction complexity increases, users tend to experience greater fatigue. As a result, CAPTCHAs such as nine-grid clicks and dynamic graphics rank lower in user preference. In contrast, color recognition and arithmetic CAPTCHAs, which are easier to operate, are favored for their simplicity but rank lowest in terms of security due to their weak protection capabilities.
We find that the public tends to prefer CAPTCHA mechanisms that are intuitive and natural in their interaction. Behavioral and shape-based CAPTCHAs are generally perceived as offering a balanced compromise between robust security and user-friendly design.
B.3 Application Trends and User Acceptance Analysis of Mainstream CAPTCHA Types
This study presents a comprehensive comparison of mainstream CAPTCHA types in terms of industry penetration, user preferences, cross-age acceptability, future growth potential, and cross-platform consistency, as summarized in Table 6. Overall, slider puzzle CAPTCHAs strike the best balance between security and usability, demonstrating broad device compatibility and promising development prospects. While image recognition and nine-grid click CAPTCHAs perform well in specific contexts, their reliance on graphic rendering and sensitivity to screen size limit their user-friendliness and scalability. Math-based and color recognition CAPTCHAs are better suited to older user groups and platforms with low graphical complexity or text-driven interfaces. Distorted and traditional text CAPTCHAs, owing to their high compatibility, continue to serve in scenarios with stringent environment adaptability requirements.
Table 6: Comparison of user experience and security performance across mainstream CAPTCHA types. This table evaluates key CAPTCHA variants based on industry adoption, user acceptance across demographics, future growth potential, and platform consistency, providing insights into usability and robustness.
CAPTCHA Type Industry Penetration (2024) Preference (Young Users) Acceptance (Middle-aged/Elderly) Growth Potential (%) Cross-Platform Consistency
Slider Puzzle CAPTCHA 68% 52% 38% 80% High (Unified on Mobile/PC)
Shape Recognition CAPTCHA 10% 12% 6% 60% Medium (Requires Graphic Rendering)
Nine-grid Click CAPTCHA 35% 28% 25% 40% Medium (Better on PC)
Mathematical Calculation CAPTCHA 12% 8% 15% 20% High (Text-Based Compatibility)
Color Recognition CAPTCHA 20% 15% 30% 60% High (Simple Graphics Adaptability)
Distorted Character CAPTCHA 5% 2% 5% 20% Medium (Font Compatibility Issues)
Traditional Character CAPTCHA 3% 1% 3% 20% Highest (Static Text Consistency)
B.4 Evaluation and Analysis of Mainstream CAPTCHA Types Selection
To comprehensively evaluate the applicability of different CAPTCHA types in real-world scenarios, this study presents a "Website CAPTCHA Selection Decision Table," constructed based on six core dimensions: security requirements, user experience, development cost, device compatibility, accessibility support, and anti-cracking update frequency. Five mainstream CAPTCHA types are analyzed and quantitatively scored using a five-point scale, where a higher score indicates better performance in the corresponding dimension, as shown in Table 7.
The sliding puzzle CAPTCHA stands out as the preferred choice for mainstream applications due to its high security, user-friendly interactive design, and strong cross-device compatibility. The nine-grid click CAPTCHA performs well in terms of interactivity and engagement but is somewhat limited by image loading speed and resolution, and shows weaknesses in accessibility support. The color recognition CAPTCHA is simple to implement, cost-effective, and highly adaptable, making it suitable for elderly users, though it lacks robustness against automated attacks. The math-based CAPTCHA offers easy deployment and strong compatibility, making it ideal for streamlined interfaces such as those in educational or government services; however, its lack of graphical interference mechanisms results in lower security. The traditional text-based CAPTCHA, while the most economical and widely supported in terms of display, is gradually being phased out due to its relatively poor performance in both security and user experience.
B.5 Analysis of CAPTCHA Error-Proneness and User Self-Recovery
This study conducts a comparative analysis of the human-computer interaction performance of six common CAPTCHA types, evaluating their error-proneness and users’ ability to self-correct mistakes in real-world scenarios. The aim is to provide data-driven insights and optimization recommendations for CAPTCHA design. Table 8 presents the average user error rate, user self-recovery rate, and common causes of errors for each CAPTCHA type.
The sliding puzzle CAPTCHA is the most user-friendly option, featuring intuitive operation and smooth interaction that enables users to quickly identify and correct errors. The nine-grid click CAPTCHA, while offering a high level of engagement, suffers from higher error rates due to small clickable areas and visual recognition challenges. The color recognition CAPTCHA introduces a novel visual design but is less accessible for users with color vision deficiencies or visual fatigue. The math-based CAPTCHA imposes a higher cognitive load, making error recovery more difficult for users. The traditional text-based CAPTCHA is hindered by distortion and blurring effects, leading to recognition difficulties and limited interaction feedback. Lastly, the shape-matching CAPTCHA relies on spatial reasoning, which may not be suitable for all users due to its higher adaptation threshold.
Table 7: Evaluation-based selection matrix for website CAPTCHA deployment. This table compares various CAPTCHA types across multiple evaluation dimensions—including security, user experience, development cost, device compatibility, accessibility, and maintenance needs—to support informed decision-making for web application integration.
Evaluation Dimension Business Requirement Priority Slider Puzzle Nine-grid Click Color Recognition Calculation Problem Traditional Character
Security Requirement High 100% 60% 40% 40% 20%
User Experience Requirement High 80% 60% 100% 40% 40%
Development Cost Low 40% 60% 80% 100% 100%
Device Compatibility Multi-device Adaptation 80% 60% 100% 100% 100%
Accessibility Support Disabled User Compatibility 60% 40% 80% 60% 40%
Anti-Cracking Tech Update Frequency High 80% 60% 40% 40% 20%
Table 8: Comparative analysis of user error and self-repair rates across CAPTCHA types. This table highlights typical failure modes encountered by users, the average error rates per CAPTCHA type, and the proportion of users able to self-correct without external assistance, providing insights into usability and design robustness.
CAPTCHA Type Average Error Rate User Self- Repair Rate Error Cause Analysis
Slider Puzzle CAPTCHA 8% 75% Inaccurate mobile touch operations (misclicking the slider area), delayed dynamic tile loading causing operation lag
Nine-grid Click CAPTCHA 15% 55% Low visibility of target graphics (confusion between similar elements), too-small click areas on mobile causing misclicks
Color Recognition CAPTCHA 12% 68% Color blindness/weakness (e.g., red-green confusion), visual fatigue from dynamic color changes
Mathematical Calculation CAPTCHA 22% 40% Misunderstanding of operation order (e.g., priority in mixed operations), ambiguous wording
Traditional Character CAPTCHA 25% 32% Excessive character distortion (blurred fonts), confusion between cases/symbols (e.g., O vs 0)
Graphic Shape Recognition CAPTCHA 18% 45% Abstract pattern comprehension difficulty (e.g., incomplete shapes), visual interference from complex background textures
B.6 Semiotic Analysis of CAPTCHA Types
From a semiotic perspective, this section presents a systematic analysis of the symbolic dimensions of various CAPTCHA types, aiming to uncover their mechanisms of meaning-making, cognitive adaptability, and symbolic evolution within human-computer interaction. Table 9 compares four major types of CAPTCHAs across several dimensions, including semantic transparency, the relationship between signifier and signified, degree of cognitive schema alignment, anti-symbolic structural capacity, and their respective stages in symbolic evolution. This analysis sheds light on the ongoing transition of CAPTCHAs from static linguistic symbols to dynamic behavioral symbols.
Text-based CAPTCHAs exhibit the highest levels of semantic transparency and signifier-signified correspondence, with a moderate cognitive load. However, their structurally regular nature makes them vulnerable to OCR-based attacks. Image-based CAPTCHAs, benefiting from intuitive visual mappings, demonstrate high schema compatibility across multilingual contexts, achieving a trade-off between usability and security. Behavioral CAPTCHAs (such as sliding puzzles and nine-grid clicks), while lower in semantic transparency and symbolic clarity, require users to quickly learn interaction rules. Nevertheless, their reliance on dynamic trajectories and spatiotemporal uncertainty grants them significant advantages in resisting symbolic deconstruction. Notably, the nine-grid click CAPTCHA—combining visual targeting with spatial interaction—is considered a "2.5D" hybrid form.
Table 9: Comparative semiotic analysis of CAPTCHA types. This table evaluates CAPTCHAs across key semiotic dimensions, including semantic transparency, signifier-signified relationships, cognitive schema alignment, and resistance to automated sign decoding, offering insight into their interpretability and robustness.
CAPTCHA Type Semantic Transparency Signifier-Signified Relevance Cognitive Schema Match Anti-sign Deconstruction Ability Stage of Sign Evolution
Traditional Character CAPTCHA High (direct text meaning) Strong Medium (language-dependent) Weak (vulnerable to OCR parsing) Linguistic signs (1D)
Graphic CAPTCHA Medium (representational images) Medium High (cross-lingual universality) Medium (AI-recognizable features) Visual signs (2D)
Behavioral CAPTCHA (Slider/Nine-grid) Low (abstract actions) Weak Low (requires rule-learning) Strong (unpredictable dynamic trajectories) Behavioral signs (3D)
Nine-grid Click CAPTCHA Medium (spatial selection) Medium Medium (game-like interaction) Medium-high (spatial ambiguity defense) Hybrid visual-behavioral signs (2.5D)
Appendix C More Extensive Experiments
C.1 Comparison of Model Stability and Human Performance Variability
The stability of model outputs across diverse task scenarios is critical to the robustness and controllability of complex CAPTCHA systems. In contrast, human participants’ performance is more susceptible to extraneous factors—such as fluctuations in attention and operator fatigue—resulting in measurable variability over repeated trials. Accordingly, this study collected multiple executions from both models and human subjects across four task categories—point-and-click recognition, static visual identification, textual logical reasoning, and interactive operations—and compared their results. As illustrated in Figure 7, boxplots and scatter plots distinctly highlight the disparity in consistency between the two, providing empirical support for subsequent stability assessment and risk management during model deployment.
Figure 7: Comparison of Multi-round Task Performance Between Models and Humans Across Task Types Each subplot illustrates a specific task type, with the left side displaying the distribution of multiple performances by human participants (scatter plots), and the right side showing the model’s stable outputs (box plots). It can be observed that the model exhibits a smaller range of variability in most tasks, indicating higher execution stability.
C.2 Error-Correction Capability Analysis
In real-world deployments, when recognition errors are inevitable, the system’s self-correction mechanism becomes a critical metric of model robustness. This study establishes a multi-stage feedback experimental framework to systematically evaluate the probability of successful iterative correction following an initial misclassification. By conducting a side-by-side comparison of model and human performance on secondary judgment tasks, we quantify the performance limits and optimization potential of existing error-correction strategies, as illustrated in Figure 8.
Figure 8: Comparison of Error Correction Capabilities Between Models and Humans Across Tasks The horizontal axis presents 20 types of interactive tasks (including representative scenarios such as Point and Click Location and State VR), while the vertical axis indicates the normalized correction capability index (ranging from 0 to 1.2). The light yellow bars represent the model’s error correction rate achieved through a multi-step feedback mechanism, whereas the ochre bars denote human participants’ spontaneous correction performance under the same tasks. The data show that in 75% of the task scenarios, the human correction index exceeds 0.8, while the model’s performance is mainly concentrated in the 0.5–0.7 range. Notably, the gap is most pronounced in TL-Q&A (dialogue-based interaction) and IM (immediate feedback) tasks. Error bars indicate the standard deviation across three independent experiments.
C.3 Comparative Analysis of Micro-Decision Path Interpretability
The interpretability of a model’s decision-making on complex multimodal CAPTCHAs is fundamental to assessing its trustworthiness in deployment. Leveraging real-world sampling data and a two-dimensional metric of mean versus variance, this study constructs a quantitative comparison framework to reveal systematic differences between model and human micro-decision processes. By concurrently analyzing decision consistency (mean) and fluctuation range (variance), we not only deconstruct the transparency of the model’s “reasoning logic” but also furnish empirical insights for enhancing its decision-making robustness, as shown in Figure 9.
Figure 9: Quantitative Comparison of Model Decision Paths in Multimodal CAPTCHA Scenarios. The horizontal axis includes 20 representative model architectures (e.g., 3-3G, HPR, J-S CC), with the left vertical axis indicating the normalized decision consistency index (ranging from 0 to 1.2), and the right vertical axis showing the variance fluctuation coefficient (ranging from 0 to 0.3). The red line represents the alignment of each model’s decision path with the human baseline, while the cobalt blue line reflects the degree of decision variability. Error bars denote standard deviations across three independent samples. The HPR model (Human-Pattern Recognition) demonstrates near-human performance in both consistency index (1.05 $\pm$ 0.07) and variance coefficient (0.12 $\pm$ 0.03). In contrast, the CD model (Cascade Decision) exhibits the highest decision variance (0.28 $\pm$ 0.05), revealing instability in decision logic. Notably, in the CCC (Cross-Channel Correlation) task, over 70% of models exceed the 0.2 threshold in variance coefficient.
Appendix D Limitations and Future Work
D.1 Limitations
Although this paper establishes the unified multimodal CAPTCHA evaluation benchmark MCA-Bench and conducts systematic testing on a LoRA-fine-tuned proxy of a visual-language model (VLM), several limitations remain. First, owing to computational constraints, our experiments focus exclusively on the QwenVL-2.5-7B-Instruct architecture. This single-model choice restricts our ability to systematically compare performance across different VLM architectures and parameter scales, and it may underrepresent the true bounds of VLM-based CAPTCHA-breaking capabilities.
Additionally, our current work emphasizes the offensive perspective—namely, the cracking performance of fine-tuned models—while defensive mechanisms receive insufficient attention. Key questions such as CAPTCHA resilience against adversarial perturbations, forged inputs, and multi-device deployment scenarios have not been explored in depth. Future studies must close the loop on the attack-defense ecology to elucidate the evolutionary dynamics of CAPTCHA systems under adversarial pressure. Moreover, although MCA-Bench covers four primary task categories (static recognition, image-click, interactive manipulation, and logical reasoning), it remains centered on mainstream CAPTCHA paradigms and lacks coverage of emerging verification scenarios, limiting its fidelity to the complexity of real-world human-machine challenges.
Finally, our training regime relies on a homogenously annotated dataset for supervised learning. In actual deployment, distribution shifts caused by device heterogeneity, varied user behaviors, and evolving task distributions may degrade model generalization and security. This issue is particularly pronounced in interactive tasks, where our simulated user trajectories—generated according to preset sampling rules—fail to capture the full spectrum of individual variability and temporal uncertainty. Consequently, the robustness of automated attacks that emulate real user behaviors may be underestimated. Enhancing the realism of data collection and the granularity of behavior modeling is therefore an essential direction for improving the benchmark’s reliability.
D.2 Future Work Outlook
MCA-Bench lays the groundwork for unified evaluation in multimodal CAPTCHA research, but its potential extends beyond attack benchmarking to the design of next-generation cognitive-security verification mechanisms. Guided by the three cognitive security design principles proposed herein—deep modality coupling, behavior-anchored verification, and semantic personalization embedding—future research should transition from “attack evaluation” toward “defensive design.” Specifically, building on modality coupling, one can devise multi-stage, multi-channel CAPTCHA workflows that intertwine visual recognition, logical reasoning, and physical interaction, thereby thwarting automation through single-path exploitation. By introducing dynamic challenge generation and adaptive difficulty adjustment based on individual user histories, systems can calibrate task complexity in real time to preserve usability while bolstering overall resistance to automated threats.